# Node Duplication Improves Cold-start Link Prediction

### v1

@@ D.8 Performance on Large-scale datasets
D.9 Ablation study
Performance with various encoders and decoders. For the ablation study, we further explored various encoders and decoders on the remaining datasets. The results are shown in Table 12 and Table 13. From these two tables, we can observe that regardless of the encoders or decoders, both NodeDup and NodeDup(L) consistently outperform the base model for Isolated and Low-degree nodes, which further demonstrates the effectiveness of our methods on cold nodes. Furthermore, NodeDup(L) consistently achieves better performance compared to the base model for Warm nodes.
@@ D.9 Ablation study
Performance with SEAL (Zhang & Chen, 2018). Considering our methods are flexible to integrate with GNN-based link prediction structures, we conduct the experiments on top of SEAL (Zhang & Chen, 2018) on the Cora and Citeseer datasets. The results are shown in Table 14. We can observe that adding NodeDup on top of SEAL can consistently improve link prediction performance in the Isolated and Low-degree node settings on these two datasets.
The influence of duplication times and nodes. In our experiments, we duplicate cold nodes once and add one edge for each cold node in NodeDup. In Figure 8, we present the results of our ablation study, focusing on the effects of duplication times and duplicated nodes on the performance of NodeDup in terms of Isolated, Low-degree, and Overall settings. The numbers displayed in each block represent the differences compared to duplicating cold nodes once. We observe that increasing the duplication times does not necessarily lead to improvements across all settings, except when duplicating all nodes for Isolated nodes performance. We also notice that duplicating all nodes multiple times can significantly enhance the performance on Isolated nodes. However, this strategy negatively impacts the overall performance due to the increased number of isolated nodes in the graph. Furthermore, we observe that duplicating cold nodes once remains the optimal strategy since it consistently delivers strong performance across all settings.
To make our analysis more comprehensive, we further conducted the experiments to show the results with duplicating warm nodes, mid-warm nodes, and randomly sampled nodes for one time, respectively, on Citeseer. The results are shown in Table 15. From the table, we can observe that duplicating mid-warm and warm nodes are not useful for the LP performance for all the settings. It’s probably because for the mid-warm and warm nodes, the neighbors’ information and supervised training signals are informative enough, therefore NodeDup cannot contribute more. We can also observe that duplicating random nodes is more effective than duplicating warm nodes but less effective than duplicating cold nodes and duplicating all nodes.
(a) Isolated
(b) Low-degree
(c) Overall
Figure 8: Ablation study on duplication times and nodes of NodeDup. (a), (b), (c) show the performance for Isolated nodes, Low-degree nodes, and Overall settings, respectively.
Table 15: Link prediction performance with different duplication nodes of NodeDup on Citeseer. "D_*" indicates duplication of "*" group nodes for one time.
Isolated Low-degree Warm Overall
Supervised 47.13 61.88 71.45 63.77
D_Isolated 54.04 72.28 74.53 69.95
D_Cold 57.54 75.50 74.68 71.73
D_Mid-warm 46.93 61.34 71.84 63.75
D_Warm 47.49 62.20 71.54 63.99
D_Random 54.10 72.39 75.05 70.06
D_All 58.87 76.09 76.01 72.44
Appendix E Implementation Details
In this section, we introduce the implementation details of our experiments. Our implementation can be found at https://github.com/snap-research/NodeDup/tree/main.
@@ Appendix E Implementation Details
Appendix F Limitations
In our work, NodeDup and NodeDup(L) are specifically proposed for LP tasks. Although cold-start is a widespread issue in all graph learning tasks, our proposed methods might not be able to generalize to other tasks, such as node classification, due to their unique design.

### v2

@@ C.10 Additional results under the inductive setting
Table 17: Performance with GCN (Kipf & Welling, 2016a) and GT (Dwivedi & Bresson, 2020) encoders, where the inner product is the decoder. The best result for each encoder is bold.
GCN GCN+NodeDup(L) GCN+NodeDup GT GT+NodeDup(L) GT+NodeDup
Cora Isolated 40.61$\pm$3.52 42.93$\pm$2.68 46.71$\pm$1.53 20.93$\pm$2.46 38.82$\pm$1.27 37.40$\pm$1.53
Low-degree 63.86$\pm$0.78 64.63$\pm$1.60 64.10$\pm$1.37 58.59$\pm$0.29 61.16$\pm$1.08 61.39$\pm$0.89
Warm 60.59$\pm$0.62 61.31$\pm$0.43 60.26$\pm$0.70 58.14$\pm$1.15 59.29$\pm$0.84 59.07$\pm$0.05
Overall 60.16$\pm$0.44 61.02$\pm$0.61 59.90$\pm$0.89 55.40$\pm$0.43 58.34$\pm$0.19 58.18$\pm$0.42
Citeseer Isolated 45.56$\pm$1.30 47.84$\pm$0.94 50.64$\pm$1.10 36.84$\pm$3.26 51.46$\pm$1.27 52.34$\pm$1.46
Low-degree 69.37$\pm$0.36 70.15$\pm$1.56 71.13$\pm$0.64 60.24$\pm$1.18 72.98$\pm$1.54 73.77$\pm$1.03
Warm 74.68$\pm$0.38 73.26$\pm$0.97 72.93$\pm$0.78 71.14$\pm$1.47 74.48$\pm$1.08 75.08$\pm$0.63
Overall 67.48$\pm$0.42 67.47$\pm$0.83 67.67$\pm$0.66 61.15$\pm$1.57 69.67$\pm$1.10 70.38$\pm$0.86
C.11 Ablation study
C.11.1 Performance with various encoders and decoders
For the ablation study, we further explored various encoders and decoders on the remaining datasets. The results are shown in Table˜15 and Table˜16. From these two tables, we can observe that regardless of the encoders or decoders, both NodeDup and NodeDup(L) consistently outperform the base model for Isolated and Low-degree nodes, which further demonstrates the effectiveness of our methods on cold nodes. Furthermore, NodeDup(L) consistently achieves better performance compared to the base model for Warm nodes.
Besides GSage, GAT and JKNet, we also conducted further experiments with convolutional-based GNNs, such as GCN (Kipf & Welling, 2016a) and GT(GraphTransformer) (Dwivedi & Bresson, 2020). The results are shown in Table˜17. Our findings indicate that our methods can also improve performance when using GCN and GT as the encoder. However, since GCN uses the same matrix for both self-representations and neighbor representations, our methods only benefit from the supervision aspect. This leads to less pronounced performance improvements on cold nodes compared to using GT and GSage as the encoder. Specifically, NodeDup shows a 13.10% improvement for GCN, 60.38% for GT, and 29.79% for GSage on isolated nodes. Moverover, NodeDup(L) on average improves GCN by 5.4%, GT by 62.58%, and GSage by 17.4%.
@@ C.11.1 Performance with various encoders and decoders
C.11.2 Performance with SEAL (Zhang & Chen, 2018)
Considering our methods are flexible to integrate with GNN-based link prediction structures, we conduct the experiments on top of SEAL (Zhang & Chen, 2018) on the Cora and Citeseer datasets. The results are shown in Table˜18. We can observe that adding NodeDup on top of SEAL can consistently improve link prediction performance in the Isolated and Low-degree node settings on these two datasets. This further demonstrates the broad applicability of NodeDup in enhancing the performance of diverse GNN-based link prediction models.
Appendix D Implementation Details
In this section, we introduce the implementation details of our experiments. Our implementation can be found at https://github.com/zhichunguo/NodeDup.
@@ Appendix D Implementation Details
Appendix E Limitations
In our work, NodeDup and NodeDup(L) are specifically proposed for LP tasks. Although cold-start is a widespread issue in all graph learning tasks, our proposed methods might not be able to generalize to other tasks, such as node classification, due to their unique design. Furthermore, the two heterophily datasets we used for evaluation involve graphs where nodes with similar features are assigned different labels. Our methods may struggle on heterophilic graphs where connected nodes have dissimilar features, such as molecular networks, which are beyond the scope of this study.
Appendix F Ethics Statement
In this work, our simple but effective method enhances the link prediction performance on cold-start nodes, which mitigates the degree bias and advances the fairness of graph machine learning. It can be widely used and beneficial for various real-world applications, such as recommendation systems, social network analysis, and bioinformatics. We do not foresee any negative societal impact or ethical concerns posed by our method. Nonetheless, we note that both positive and negative societal impacts can be made by applications of graph machine learning techniques, which may benefit from the improvements induced by our work. Care must be taken, in general, to ensure positive societal and ethical consequences of machine learning.